OpenStudy (anonymous):

For the function F(x)=(1/3)x^3 +x^2 -3x +6, find the coordinates of the points with zero gradients and sketch the graph of F(x), showing the turning points and y-intercept

OpenStudy (jack1):

so \[\large F(x)=\frac 13 x^3 +x^2 -3x +6\] points with zero gradients: first find derivative of F(x) make F'(x) = 0 sub those x coordinates found back into F(x) to get y coordinates =turning points of function

OpenStudy (jack1):

y intercept can be read straight off original function \[F(x)=(1/3)x^3 +x^2 -3x +6\] \[F(x)= a x^3 + bx^2 -cx +d\] y intercept is "d"

OpenStudy (anonymous):

thankyou!